OpenStudy (anonymous):

When solving x2 – 16x = 10 by completing the square what value is added to both sides of the equation? 6 8 16 64

jhonyy9 (jhonyy9):

- so if we considering (x-8) squared so than this is equal x2 -16x +64 - so x2 -16x=10 what is x2 -16x -10=0 - so added to both sides anything from these not is possible making complete square